a b 5. (10 points) Find all matrices A= с d ſi 1 that satisfy A (3-61 A. 1 1
Answer #1

s) A= Foel A [edit] 03-07 :16] [at [a led C d atb ath Cat ata b+d brd Ctd ced ate atb=a+C > b=c atbaitd =) a=d C+d=ate =) d=a